OpenStudy (anonymous):

Fill in the missing exponents (3^9*3^3)/(3^8*3^2)=3

OpenStudy (anonymous):

(9+ 3) - (8 + 2) = x, solve for x

OpenStudy (anonymous):

(3^9*3^3)/(3^8*3^2) = 3^2

OpenStudy (anonymous):

9...

OpenStudy (anonymous):

9...